OpenStudy (anonymous):

The perimeter of a rectangular garden is 54yd. One side has a length of 15yd. Find the area of the garden. What do you need to know to find the area of the garden? How can you use the perimeter and one side to find the area? What is the length of the other side of the garden?

OpenStudy (ddcamp):

P = 2l + 2w l = 15 54 = 2(15) + 2w 54 = 30 + 2w 24 = 2w 12 = w A = l*w = 12*15 = 180